Plasma is the fluid portion of blood. The concentratio of asprin ( molar mass= 180.0g/mol) in your plasma is found to be 2.99x 10-4 M after you take 2 tablets of asprin. If the volume of your plasma is 5.85 L, how many grams of aspirin are in your blood? ( assume density of plasma is 1.00g/mol.)
1
Expert's answer
2012-02-17T08:33:03-0500
V=5.85 L
C=2.99x 10-4 M

2.99x10-4 moles in 1 litre
X moles & & & & & in
5.85 L

X=(2.99x10-4* 5.85)/ 1= 1.749x10-3
m=n*Mm=1.749x10-3
* 180.0=0.3148
m=0,3148 g